OpenStudy (anonymous):

What is the height of a triangle that has an area of 60 yd2 and a base with a length of 12 yd?

OpenStudy (anonymous):

area = 1/2 (b*h) sub in values 60 = (b*12)/2 solve 120 = 12b 10 = b base length = 10 yd